OpenStudy (goldrush18):

the circle C1 has (-3,4) and (1,2) as endpoints of a diameter. i. Show that the equation of C1 is x^2+y^2+2x-6y+5=0. ii. The Circle C2 has equation x^2+y^2+x-5y=0. Calculate the coordinates of the point of intersection of C1 and C2

OpenStudy (anonymous):

since the 2 points are the end points of diameter center is given by (-3+1)/2 ,(4+2)/2=(-1,3)..and radius = sqrt ((-1+3)^2+(3-4)^2)=sqrt 5..so circle equation is (x+1)^2 +(y-3)^2=5 simplifying x^2+y^2+2x-6y+5=0.

OpenStudy (goldrush18):

ok thanks